jaylaa1:

To attend the football game, it costs $8 for parking plus $59 per ticket. If Bryan paid #362, how many tickets did he purchase?

jaylaa1:

help please

jaylaa1:

$362^

arrivhs:

59x+8= 362 59x= 362-8 59x= 354 X= 354/59 X= 6

arrivhs:

the answer is 6

TETSXPREME:

wait

jaylaa1:

wait so is the answer 6?

arrivhs:

@jaylaa1 wrote:
wait so is the answer 6?
yes

jaylaa1:

cause i would do 59x+8= 362 59x= 362-8 59x= 354 X= 354/59 X= 6 like this right?

arrivhs:

@jaylaa1 wrote:
cause i would do 59x+8= 362 59x= 362-8 59x= 354 X= 354/59 X= 6 like this right?
yes

jaylaa1:

okay thanks

arrivhs:

@jaylaa1 wrote:
okay thanks
your welcome

Brunoooooooo:

Yeah Arrivhs is doing it right

RodWav3:

you dont knoww this?

RodWav3:

I TAUGHT U BETTER

jaylaa1:

shush mar

jaylaa1:

IMAO

Brunoooooooo:

lmao

arrivhs:

im in 8th grade

RodWav3:

IG u need more lessons yk

jaylaa1:

i feel ashamed now IMAO

Brunoooooooo:

😂

RodWav3:

@jaylaa1 wrote:
i feel ashamed now IMAO
Lol

jaylaa1:

they really just called me out lmaoo

Brunoooooooo:

Dayum o-o

RodWav3:

@jaylaa1 wrote:
they really just called me out lmaoo
sorry ...

RodWav3:

@rodwav3 wrote:
@jaylaa1 wrote:
they really just called me out lmaoo
sorry ...
not sorry

arrivhs:

Im soooo s-sad

jaylaa1:

IMAOO

RodWav3:

@arrivhs wrote:
Im soooo s-sad
dont be

arrivhs:

My feelings os awalys showing and im scared

jaylaa1:

wait ur being sarcastic right?? ;-;

jaylaa1:

im sorry

arrivhs:

Im going to gp kms Byeeeeeeeeee

RodWav3:

@jaylaa1 wrote:
im sorry
....

jaylaa1:

AYOOO

jaylaa1:

WAITTT

RodWav3:

@arrivhs wrote:
Im going to gp kms Byeeeeeeeeee
CHILLLLLLL

jaylaa1:

THIS ESCALATED QUICKLY

arrivhs:

No i can die if i want

jaylaa1:

@rodwav3 they toldddd u lol

RodWav3:

@jaylaa1 wrote:
@rodwav3 they toldddd u lol
told youu